if cos theta = -5/13 and sin theta > 0 , then tan theta is A)-12/5 B)-5/12 C)5/12 D)12/5
Answer by amarjeeth123(570)

cos theta = -5/13
sin theta > 0
sin theta=12/13
tan theta=sin theta/cos theta=-12/5.